OpenStudy (anonymous):

determine all the values of c so funtion is continuous on (-infinity, infinity)

OpenStudy (anonymous):

f(x) + x^2 x<c and x+1 x> or equal to c

OpenStudy (anonymous):

at x=c x^2= x+1 x^2-x-1= 0 then c= (1+or- sqrt(5))/2